Presentation is loading. Please wait.

Presentation is loading. Please wait.

TEC X5 Setup (1) 1 Front + 1 Back petal, fully equipped (28 modules F.P. + 23 modules B.P. + AOH + DOHM) Petal assembly IPN-Lyon (F.P.), RWTH-Aachen I.

Similar presentations


Presentation on theme: "TEC X5 Setup (1) 1 Front + 1 Back petal, fully equipped (28 modules F.P. + 23 modules B.P. + AOH + DOHM) Petal assembly IPN-Lyon (F.P.), RWTH-Aachen I."— Presentation transcript:

1 TEC X5 Setup (1) 1 Front + 1 Back petal, fully equipped (28 modules F.P modules B.P. + AOH + DOHM) Petal assembly IPN-Lyon (F.P.), RWTH-Aachen I (B.P.) Constitutes one full TEC control ring Full optical link with optical PMC-FEC (VME- FEC will be added in coming days) 2 x FED9U in same crate (68 chans F.P., 56 chans B.P.)

2 TEC X5 Setup (2) Petals placed back-to-back in insulated cold / dark box Series cooling loop: B.P. -> F.P. -> copper cooling plate Cooling plant C Temp. / Humidity monitoring + DCU information being read

3 Installation experience Mechanical installation, cabling, etc.: 2 days FEC communication established immediately First Landau: 4 days after arrival Approximate strip statistics after installation at X5 - 1 dead readout channel (256 strips) not included: – Dead: 99 / > 0.31 % – Noisy: 8 / > 0.03 % –(No loss of strips after 800 km car transport Aachen to CERN) One module with HV short

4 Early Results FED9U works – Internal pedestal + common mode noise subtraction – Zero suppression being studied – first results encouraging ENC noise close to expected: –Estimated independently using (a) calibration pulses (b) signal peak (assume 22,000 e- in 320 micron Si) –Fit with known detector capacitance gives: ENC = C (C ranges 11 pF R1 to 26 pF R7)

5 ENC noise from cal runs Note: online data, very rough analysis!

6 ENC noise from signal peak

7 Voltage scan – FP R1/R2

8 Voltage scan – BP R1/R2

9 Selected Landaus, Peak Mode V = 300 V

10 S/N vs Ring (Peak Mode) V=300 V

11 FED Zero Suppression Note: Problem with internal FED common mode subtraction for these runs

12 First results at low temp. Temperature of sensor leg = -5 °C approx. (Compare S/N room temp = 18)

13 Problems One dead R5 readout channel (256 strips) HV short on FP stereo module 2.4 BP 2.1 does not receive HV One R5 module with very low noise (broken bonds?) High CMN on R5 module associated with noisy strip

14 F.P. 2.1 does not get HV

15 Dead readout channel

16 ???

17 ST sensor CMN problem

18 To Do this Week No beam Improve thermal contact of T and H sensors on modules Improve cooling circuit and thermal isolation of box Logging of DCU values Commission VME FEC Try higher fast control speeds Pulse shape studies (cal runs) Repeat bad strip study Finalize installation of slow ctrl / interlocks


Download ppt "TEC X5 Setup (1) 1 Front + 1 Back petal, fully equipped (28 modules F.P. + 23 modules B.P. + AOH + DOHM) Petal assembly IPN-Lyon (F.P.), RWTH-Aachen I."

Similar presentations


Ads by Google